# Exact Exploratory Bi-factor Analysis: A Constraint-based Optimisation Approach

### v1

@@ A.1 Constrained Optimisation for Exploratory Hierarchical Factor Analysis
Figure A.1: The illustrative example of a three-layer hierarchical factor model.
Panel (a) of Figure A.1 provides the path diagram of a hierarchical factor model that has three layers, with factor $F_{1}$ in layer 3, factors $F_{2}$ and $F_{3}$ in layer 2, and factors $F_{4}$-$F_{7}$ in layer 1. The corresponding factor hierarchy is summarised in Panel (b) of Figure A.1 that takes the form of a tree, where $F_{2}$ and $F_{3}$ are nested within $F_{1}$, $F_{4}$ and $F_{5}$ are nested within $F_{2}$, and $F_{6}$ and $F_{7}$ are nested within $F_{3}$. In what follows, we show how the loading structure of this three-layer hierarchical model can be learned by a constrained optimisation method, assuming that the factor hierarchy in Panel (b) of Figure A.1 is known while the variables loading on each factor are unknown. The goal is to learn how the observed variables load on the seven factors.
@@
$\lambda_{j2}\lambda_{j3}=0,~{}~{}\lambda_{j2}\lambda_{j6}=0,~{}~{}\lambda_{j2}\lambda_{j7}=0,$ (A.1)
@@
Consequently, the corresponding hierarchical factor model can be learned by minimising the loss function $l(\Lambda\Lambda^{\top}+\Psi(\bm{\psi});S)$, subject to the constraints in (A.1).
Although the above discussion focuses on the specific hierarchical factor structure in Figure A.1, when given a different factor hierarchy, it is easy to derive similar constraints as in (A.1) by induction. Based on the constraints, the corresponding hierarchical factor model can be learned by an ALM.
Finally, we note that the factor hierarchy is typically unknown in practice. In that case, we need an algorithm that simultaneously learns the factor hierarchy and the variable loadings on the hierarchical factors. As there are exponentially many choices for the structure of factor hierarchy, this problem is more challenging than the setting when the factor hierarchy is known. It is also more challenging than exploratory bi-factor analysis with unknown group factors, as the bi-factor model has a simple two-layer factor hierarchy that is completely determined by the number of factors. See further discussions on this learning problem in Section 5.
A.2 Simulation
In this section, we examine the recovery of the hierarchical structure of our method. For $J\in\{20,40\}$ and $N\in\{500,2000\}$, a data generation model is considered, resulting in a total 3 simulation settings. With slight abuse of notation, we denote by $\mathcal{B}_{g}^{*}$ as the true item groups related with the $g$th factor. In the data generation model, $B_{1}^{*}=\{1,\ldots,J\}$, $B_{2}^{*}=\{1,\ldots,J/2\}$, $B_{3}^{*}=\{J/2,\ldots,J\}$, $B_{4}^{*}=\{1,\ldots,J/4\}$, $B_{5}^{*}=\{J/4,\ldots,J/2\}$, $B_{6}^{*}=\{J/2,\ldots,3J/4\}$, $B_{7}^{*}=\{3J/4,\ldots,J\}$. $\Psi^{*}=\mathbb{I}_{J\times J}$, and $\Lambda^{*}$ follows
@@ A.2 Simulation
for $j=1,\ldots,J$ and $k=1,\ldots,G+1$. Here $u_{jk}$ follows i.i.d Uniform(0,1) distribution and $x_{jk}$ follows i.i.d standard normal distribution.
@@
For each setting, we first generate $\Lambda^{*}$ once and use them to generate 100 datasets. The averaged results under 100 replication are shown in Table A.1. From the simulation results, we find that our method performs well on the recovery of hierarchical structure.
Table A.1: Simulation results of the recovery of hierarchical factor structure.
@@
20 500 0.83 0.99
2000 0.92 0.99
40 500 1.00 1.00
2000 1.00 1.00
Appendix B Extraversion Scale Item Key
Table B.1: Extraversion Item Key
@@ Appendix B Extraversion Scale Item Key
Appendix C Technical Proofs
C.1 Proof of Theorem 1
@@
$\Lambda[\mathcal{B}^{*}_{g_{1}},:]\Phi(\Lambda[\mathcal{B}^{*}_{g_{1}},:])^{\top}=\Lambda^{*}[\mathcal{B}^{*}_{g_{1}},\{1,1+g_{1}\}](\Lambda^{*}[\mathcal{B}^{*}_{g_{1}},\{1,1+g_{1}\}])^{\top}.$ (C.1)
Since the right side matrix has rank 2, there exist 2 possible bi-factor structures of the left side: ($\Lambda[\mathcal{B}^{*}_{g_{1}},\{1,1+g^{\prime}_{1}\}](\Lambda[\mathcal{B}^{*}_{g_{1}},\{1,1+g^{\prime}_{1}\}])^{\top}=\Lambda^{*}[\mathcal{B}^{*}_{g_{1}},\{1,1+g_{1}\}](\Lambda^{*}[\mathcal{B}^{*}_{g_{1}},\{1,1+g_{1}\}])^{\top}$ for some $g^{\prime}_{1}\in\{1,\ldots,G\}$and (2) There exists a partition of $\mathcal{B}^{*}_{g_{1}}=\mathcal{B}^{*}_{g_{1,1}}\cup\mathcal{B}^{*}_{g_{1,2}}$ and $g^{\prime}_{1},g^{\prime}_{2}\in\{1,\ldots,G\}$ such that
$\begin{pmatrix}\bm{\lambda}_{1}&\bm{\lambda}_{g^{\prime}_{1}}&\mathbf{0}\\ \bm{\lambda}_{2}&\mathbf{0}&\bm{\lambda}_{g^{\prime}_{2}}\end{pmatrix}\begin{pmatrix}1&0&0\\ 0&1&\phi_{1+g^{\prime}_{1},1+g^{\prime}_{2}}\\ 0&\phi_{1+g^{\prime}_{1},1+g^{\prime}_{2}}&1\end{pmatrix}\begin{pmatrix}\bm{\lambda}_{1}^{\top}&\bm{\lambda}_{2}^{\top}\\ \bm{\lambda}_{g^{\prime}_{1}}^{\top}&\mathbf{0}^{\top}\\ \mathbf{0}^{\top}&\bm{\lambda}_{g^{\prime}_{2}}^{\top}\end{pmatrix}=\begin{pmatrix}\bm{\lambda}^{*}_{1}&\bm{\lambda}^{*}_{g_{1}}\\ \bm{\lambda}^{*}_{2}&\bm{\lambda}^{*}_{g_{2}}\end{pmatrix}\begin{pmatrix}(\bm{\lambda}^{*}_{1})^{\top}&(\bm{\lambda}^{*}_{2})^{\top}\\ (\bm{\lambda}^{*}_{g_{1}})^{\top}&(\bm{\lambda}^{*}_{g_{2}})^{\top}\end{pmatrix},$ (C.2)
@@
Here we consider the second case. Since the matrix on the right side has rank 2, we must have $(\bm{\lambda}_{i},\bm{\lambda}_{g^{\prime}_{i}})$ has rank 1 for $i=1,2$, which leads to the fact that $(\bm{\lambda}^{*}_{i},\bm{\lambda}^{*}_{g_{i}})$ has rank 1. However, by Condition 2, there exists at least one of $(\bm{\lambda}^{*}_{1},\bm{\lambda}^{*}_{g_{1}})$ and $(\bm{\lambda}^{*}_{2},\bm{\lambda}^{*}_{g_{2}})$ has rank 2. Thus, we must have $\Lambda[\mathcal{B}^{*}_{g_{1}},\{1,1+g^{\prime}_{1}\}](\Lambda[\mathcal{B}^{*}_{g_{1}},\{1,1+g^{\prime}_{1}\}])^{\top}=\Lambda^{*}[\mathcal{B}^{*}_{g_{1}},\{1,1+g_{1}\}](\Lambda^{*}[\mathcal{B}^{*}_{g_{1}},\{1,1+g_{1}\}])^{\top}$ for some $g^{\prime}_{1}\in\{1,\ldots,G\}$. Without loss of generation, we assume $g^{\prime}_{1}=g_{1}$.
@@
$\begin{pmatrix}\bm{\lambda}_{1}&\bm{\lambda}_{g_{1}}&\mathbf{0}\\ \bm{\lambda}_{2}&\mathbf{0}&\bm{\lambda}_{g^{\prime}_{2}}\end{pmatrix}\begin{pmatrix}1&0&0\\ 0&1&\rho_{1,2}\\ 0&\rho_{1,2}&1\end{pmatrix}\begin{pmatrix}\bm{\lambda}_{1}^{\top}&\bm{\lambda}_{2}^{\top}\\ \bm{\lambda}_{g_{1}}^{\top}&\mathbf{0}^{\top}\\ \mathbf{0}^{\top}&\bm{\lambda}_{g^{\prime}_{2}}^{\top}\end{pmatrix}=\begin{pmatrix}\bm{\lambda}^{*}_{1}&\bm{\lambda}^{*}_{g_{1}}&\mathbf{0}\\ \bm{\lambda}^{*}_{2}&\mathbf{0}&\bm{\lambda}^{*}_{g_{2}}\end{pmatrix}\begin{pmatrix}1&0&0\\ 0&1&\rho_{1,2}^{*}\\ 0&\rho_{1,2}^{*}&1\end{pmatrix}\begin{pmatrix}(\bm{\lambda}_{1}^{*})^{\top}&(\bm{\lambda}_{2}^{*})^{\top}\\ (\bm{\lambda}_{g_{1}}^{*})^{\top}&\mathbf{0}^{\top}\\ \mathbf{0}^{\top}&(\bm{\lambda}_{g_{2}}^{*})^{\top},\end{pmatrix}$ (C.3)
@@
$\begin{pmatrix}\bm{\lambda}_{1}&\bm{\lambda}_{g_{1}}&\mathbf{0}\\ \bm{\lambda}_{2,1}&\bm{\lambda}_{2,g_{1}}&\mathbf{0}\\ \bm{\lambda}_{2,2}&\mathbf{0}&\bm{\lambda}_{2,g^{\prime}_{2}}\end{pmatrix}\begin{pmatrix}1&0&0\\ 0&1&\rho_{1,2}\\ 0&\rho_{1,2}&1\end{pmatrix}\begin{pmatrix}\bm{\lambda}_{1}^{\top}&\bm{\lambda}_{2,1}^{\top}&\bm{\lambda}_{2,2}^{\top}\\ \bm{\lambda}_{g_{1}}^{\top}&\bm{\lambda}_{2,g_{1}}^{\top}&\mathbf{0}^{\top}\\ \mathbf{0}^{\top}&\mathbf{0}^{\top}&\bm{\lambda}_{2,g^{\prime}_{2}}^{\top}\end{pmatrix}$ (C.4)

### v2

@@ D.1 Constrained Optimisation for Exploratory Hierarchical Factor Analysis
Figure D.1: The illustrative example of a three-layer hierarchical factor model.
Panel (a) of Figure D.1 provides the path diagram of a hierarchical factor model that has three layers, with factor $F_{1}$ in layer 3, factors $F_{2}$ and $F_{3}$ in layer 2, and factors $F_{4}$-$F_{7}$ in layer 1. The corresponding factor hierarchy is summarised in Panel (b) of Figure D.1 that takes the form of a tree, where $F_{2}$ and $F_{3}$ are nested within $F_{1}$, $F_{4}$ and $F_{5}$ are nested within $F_{2}$, and $F_{6}$ and $F_{7}$ are nested within $F_{3}$. In what follows, we show how the loading structure of this three-layer hierarchical model can be learned by a constrained optimisation method, assuming that the factor hierarchy in Panel (b) of Figure D.1 is known while the variables loading on each factor are unknown. The goal is to learn how the observed variables load on the seven factors.
@@
$\lambda_{j2}\lambda_{j3}=0,~{}~{}\lambda_{j2}\lambda_{j6}=0,~{}~{}\lambda_{j2}\lambda_{j7}=0,$ (D.1)
@@
Consequently, the corresponding hierarchical factor model can be learned by minimising the loss function $l(\Lambda\Lambda^{\top}+\Psi(\bm{\psi});S)$, subject to the constraints in (D.1).
Although the above discussion focuses on the specific hierarchical factor structure in Figure D.1, when given a different factor hierarchy, it is easy to derive similar constraints as in (D.1) by induction. Based on the constraints, the corresponding hierarchical factor model can be learned by an ALM.
Finally, we note that the factor hierarchy is typically unknown in practice. In that case, we need an algorithm that simultaneously learns the factor hierarchy and the variable loadings on the hierarchical factors. As there are exponentially many choices for the structure of factor hierarchy, this problem is more challenging than the setting when the factor hierarchy is known. It is also more challenging than exploratory bi-factor analysis with unknown group factors, as the bi-factor model has a simple two-layer factor hierarchy that is completely determined by the number of factors.
D.2 Simulation
In this section, we examine the recovery of the hierarchical structure of our method. For $J\in\{20,40\}$ and $N\in\{500,2000\}$, a data generation model is considered, resulting in a total of 4 simulation settings. With slight abuse of notation, we denote by $\mathcal{B}_{g}^{*}$ as the true item groups related to the $g$th factor. In the data generation model, $B_{1}^{*}=\{1,\ldots,J\}$, $B_{2}^{*}=\{1,\ldots,J/2\}$, $B_{3}^{*}=\{J/2,\ldots,J\}$, $B_{4}^{*}=\{1,\ldots,J/4\}$, $B_{5}^{*}=\{J/4,\ldots,J/2\}$, $B_{6}^{*}=\{J/2,\ldots,3J/4\}$, $B_{7}^{*}=\{3J/4,\ldots,J\}$. $\Psi^{*}=\mathbb{I}_{J\times J}$, and $\Lambda^{*}$ follows
@@ D.2 Simulation
for $j=1,\ldots,J$ and $k=1,\ldots,G+1$. Here, $u_{jk}$s are i.i.d., following a Uniform$(0.2,1)$ distribution, and $x_{jk}$s are i.i.d., following a Bernoulli$(0.5)$ distribution.
@@
For each setting, we first generate $\Lambda^{*}$ once and use them to generate 100 datasets. The averaged results under 100 replication are shown in Table D.1. From the simulation results, we find that our method performs well on the recovery of hierarchical structure.
Table D.1: Simulation results of the recovery of hierarchical factor structure.
@@
20 500 0.94 0.998
2000 1.00 1.000
40 500 0.88 0.988
2000 1.00 1.000
Appendix E Extraversion Scale Item Key
Table E.1: Extraversion Item Key
@@ Appendix E Extraversion Scale Item Key
Appendix F Real Data Analysis using Bi-factor Rotation Method
In this section, we present the results of the same data in Section 4 by bi-factor rotation method as a comparison with our proposed method.
Using a candidate set $\mathcal{G}=\{2,\ldots,12\}$, the BIC procedure of exploratory factor analysis given in Section 3.2 selects eight factors in total, which coincide with the number of factors selected by the BIC procedure of our proposed method. By applying the bi-factor rotation method(Jennrich and Bentler,, 2012), we get the rotation solutions $\widehat{\Lambda}^{oblq}$ in Table F.1 and $\widehat{\Phi}^{oblq}$ in equation (F.1).
Table F.1: Estimated loading matrix $\widehat{\Lambda}^{oblq}$ with seven group factors.
Items Sign General G1 G2 G3 G4 G5 G6 G7
1 +E1 0.86 0.01 -0.06 -0.08 -0.03 -0.04 0.42 -0.08
2 +E1 0.85 0.03 -0.11 0.03 0.06 -0.12 0.07 -0.01
3 --E1 0.91 -0.02 -0.12 -0.01 0.03 -0.09 -0.02 -0.05
4 --E1 0.87 -0.14 -0.04 -0.01 -0.10 -0.14 -0.03 -0.20
5 +E2 0.88 0.69 0.00 0.00 -0.02 0.00 -0.01 0.01
6 +E2 0.92 0.25 0.03 -0.12 0.09 0.05 0.22 -0.03
7 --E2 0.72 -0.06 -0.03 -0.04 -0.08 -0.16 -0.21 -0.12
8 --E2 0.85 0.22 -0.02 -0.05 -0.06 -0.07 -0.29 -0.08
9 +E3 0.52 0.02 0.02 0.00 0.79 0.02 0.04 -0.03
10 +E3 0.52 0.04 0.00 0.00 0.75 -0.03 0.03 0.01
11 +E3 0.44 -0.03 0.00 0.05 0.62 0.04 -0.08 0.01
12 --E3 0.55 -0.09 -0.04 -0.02 0.62 -0.05 -0.06 0.06
13 +E4 0.32 0.05 0.01 0.00 0.02 0.82 -0.02 -0.06
14 +E4 0.51 -0.07 0.02 -0.02 -0.01 0.74 0.06 0.06
15 +E4 0.49 0.02 -0.08 0.15 -0.02 0.51 -0.06 0.14
16 --E4 0.19 -0.14 -0.04 -0.14 0.08 0.37 -0.19 -0.07
17 +E5 0.46 0.09 -0.03 -0.04 0.02 0.02 0.06 0.49
18 +E5 0.53 -0.05 0.02 0.00 0.01 0.03 -0.04 0.62
19 +E5 0.28 0.05 0.48 -0.02 0.02 -0.12 0.00 0.33
20 +E5 0.48 0.00 1.10 0.00 0.00 0.01 0.00 -0.02
21 +E6 0.64 -0.12 0.04 0.26 -0.04 0.01 0.28 -0.01
22 +E6 0.69 0.06 0.11 0.40 -0.01 0.01 0.09 0.08
23 +E6 0.63 0.02 -0.01 0.63 0.03 0.01 -0.05 -0.01
24 +E6 0.58 -0.04 -0.03 0.60 0.00 -0.02 0.01 -0.04
$\widehat{\Phi}^{oblq}=\left(\begin{matrix}1&0&0&0&0&0&0&0\\ 0&1&0.19&-0.14&-0.17&-0.15&0.10&0.05\\ 0&0.19&1&-0.10&0.00&-0.07&0.11&0.39\\ 0&-0.14&-0.10&1&0.02&0.07&0.06&0.07\\ 0&-0.17&0.00&0.02&1&0.21&-0.07&0.11\\ 0&-0.15&-0.07&0.07&0.21&1&-0.09&0.01\\ 0&0.10&0.11&0.06&-0.07&-0.09&1&0.00\\ 0&0.05&0.39&0.07&0.11&0.01&0.00&1\\ \end{matrix}\right).$ (F.1)
To help to identify a bi-factor structure from $\widehat{\Lambda}^{oblq}$, all loadings whose absolute value is less than 0.2 are set to zero, as is done in Jennrich and Bentler, (2012). The adjusted loadings are presented in Table F.2. As expected, the loading structure does not conform strictly to a bi-factor model, with four items loading onto three factors.
Table F.2: Estimated bi-factor loading matrix with seven group factors.
Items Sign General G1 G2 G3 G4 G5 G6 G7
1 +E1 0.86 0 0 0 0 0 0.42 0
2 +E1 0.85 0 0 0 0 0 0 0
3 --E1 0.91 0 0 0 0 0 0 0
4 --E1 0.87 0 0 0 0 0 0 -0.20
5 +E2 0.88 0.69 0 0 0 0 0 0
6 +E2 0.92 0.25 0 0 0 0 0.22 0
7 --E2 0.72 0 0 0 0 0 -0.21 0
8 --E2 0.85 0.22 0 0 0 0 -0.29 0
9 +E3 0.52 0 0 0 0.79 0 0 0
10 +E3 0.52 0 0 0 0.75 0 0 0
11 +E3 0.44 0 0 0 0.62 0 0 0
12 --E3 0.55 0 0 0 0.62 0 0 0
13 +E4 0.32 0 0 0 0 0.82 0 0
14 +E4 0.51 0 0 0 0 0.74 0 0
15 +E4 0.49 0 0 0 0 0.51 0 0
16 --E4 0 0 0 0 0 0.37 0 0
17 +E5 0.46 0 0 0 0 0 0 0.49
18 +E5 0.53 0 0 0 0 0 0 0.62
19 +E5 0.28 0 0.48 0 0 0 0 0.33
20 +E5 0.48 0 1.10 0 0 0 0 0
21 +E6 0.64 0 0 0.26 0 0 0.28 0
22 +E6 0.69 0 0 0.40 0 0 0 0
23 +E6 0.63 0 0 0.63 0 0 0 0
24 +E6 0.58 0 0 0.60 0 0 0 0
We now analyze the estimated model in detail. In this result, we have adjusted the sign flip and column swapping to align with the result of the proposed method. All loadings on the general factor are positive, supporting the existence of a general extraversion factor. We interpret the group factors G3, G4, G5 as the Cheerfulness, Assertiveness and Activity Level factors respectively. G2, loaded with the items ”19 Enjoy being reckless” and ”20 Act wild and crazy” is interpreted as the Reckless Excitement-Seeking factor and consistent with the result from our proposed method. G7 is loaded with items ”4 Keep others at a distance”, ”17 Love excitement”, ”18 Seek adventure” and ”19 Enjoy being reckless”. Even though G2 and G7 are loaded with item 19 in common, G7 emphasizes more on the pursuit of meaningful experiences. So we still interpret G7 as the Meaningful Excitement-Seeking factor. Additionally, G2 and G7 are positively correlated, as is the case in Section 4.
There is a notable difference between the results from the two methods. The result of the ALM method shows the clear presence of a Friendliness factor (G1) and a Gregariousness factor (G6). However, for the bi-factor rotation method, these does not seem to exist a clear Friendliness factor. Both G1 and G6 in the solution of the bi-factor rotation method are related to Gregariousness. Large loadings of the variables designed to measure Friendliness now spread out among several group factors.
Overall, both methods suggest similar (approximate) bi-factor model structures, and the result from the proposed method tends to be neater and more interpretable.
Appendix G Technical Proofs
G.1 Proof of Theorem 1
@@
$\Lambda[\mathcal{B}^{*}_{g_{1}},:]\Phi(\Lambda[\mathcal{B}^{*}_{g_{1}},:])^{\top}=\Lambda^{*}[\mathcal{B}^{*}_{g_{1}},\{1,1+g_{1}\}](\Lambda^{*}[\mathcal{B}^{*}_{g_{1}},\{1,1+g_{1}\}])^{\top}.$ (G.1)
Since the matrix on the right side of (G.1) has rank 2, there exist 2 possible bi-factor structures for the matrix on the left side of (G.1): ($\Lambda[\mathcal{B}^{*}_{g_{1}},\{1,1+g^{\prime}_{1}\}](\Lambda[\mathcal{B}^{*}_{g_{1}},\{1,1+g^{\prime}_{1}\}])^{\top}=\Lambda^{*}[\mathcal{B}^{*}_{g_{1}},\{1,1+g_{1}\}](\Lambda^{*}[\mathcal{B}^{*}_{g_{1}},\{1,1+g_{1}\}])^{\top}$ for some $g^{\prime}_{1}\in\{1,\ldots,G\}$ and (2) There exists a partition of $\mathcal{B}^{*}_{g_{1}}=\mathcal{B}^{*}_{g_{1,1}}\cup\mathcal{B}^{*}_{g_{1,2}}$ and $g^{\prime}_{1},g^{\prime}_{2}\in\{1,\ldots,G\}$ such that
$\begin{pmatrix}\bm{\lambda}_{1}&\bm{\lambda}_{g^{\prime}_{1}}&\mathbf{0}\\ \bm{\lambda}_{2}&\mathbf{0}&\bm{\lambda}_{g^{\prime}_{2}}\end{pmatrix}\begin{pmatrix}1&0&0\\ 0&1&\phi_{1+g^{\prime}_{1},1+g^{\prime}_{2}}\\ 0&\phi_{1+g^{\prime}_{1},1+g^{\prime}_{2}}&1\end{pmatrix}\begin{pmatrix}\bm{\lambda}_{1}^{\top}&\bm{\lambda}_{2}^{\top}\\ \bm{\lambda}_{g^{\prime}_{1}}^{\top}&\mathbf{0}^{\top}\\ \mathbf{0}^{\top}&\bm{\lambda}_{g^{\prime}_{2}}^{\top}\end{pmatrix}=\begin{pmatrix}\bm{\lambda}^{*}_{1}&\bm{\lambda}^{*}_{g_{1}}\\ \bm{\lambda}^{*}_{2}&\bm{\lambda}^{*}_{g_{2}}\end{pmatrix}\begin{pmatrix}(\bm{\lambda}^{*}_{1})^{\top}&(\bm{\lambda}^{*}_{2})^{\top}\\ (\bm{\lambda}^{*}_{g_{1}})^{\top}&(\bm{\lambda}^{*}_{g_{2}})^{\top}\end{pmatrix},$ (G.2)
@@
Here we consider the second case. Since the matrix on the right side of (G.2) has rank 2, we must have $(\bm{\lambda}_{i},\bm{\lambda}_{g^{\prime}_{i}})$ has rank 1 for $i=1,2$, which leads to the fact that $(\bm{\lambda}^{*}_{i},\bm{\lambda}^{*}_{g_{i}})$ has rank 1. However, by Condition 2, there exists at least one of $(\bm{\lambda}^{*}_{1},\bm{\lambda}^{*}_{g_{1}})$ and $(\bm{\lambda}^{*}_{2},\bm{\lambda}^{*}_{g_{2}})$ has rank 2. Thus, we must have $\Lambda[\mathcal{B}^{*}_{g_{1}},\{1,1+g^{\prime}_{1}\}](\Lambda[\mathcal{B}^{*}_{g_{1}},\{1,1+g^{\prime}_{1}\}])^{\top}=\Lambda^{*}[\mathcal{B}^{*}_{g_{1}},\{1,1+g_{1}\}](\Lambda^{*}[\mathcal{B}^{*}_{g_{1}},\{1,1+g_{1}\}])^{\top}$ for some $g^{\prime}_{1}\in\{1,\ldots,G\}$. Without loss of generation, we assume $g^{\prime}_{1}=g_{1}$.
@@
$\begin{pmatrix}\bm{\lambda}_{1}&\bm{\lambda}_{g_{1}}&\mathbf{0}\\ \bm{\lambda}_{2}&\mathbf{0}&\bm{\lambda}_{g^{\prime}_{2}}\end{pmatrix}\begin{pmatrix}1&0&0\\ 0&1&\rho_{1,2}\\ 0&\rho_{1,2}&1\end{pmatrix}\begin{pmatrix}\bm{\lambda}_{1}^{\top}&\bm{\lambda}_{2}^{\top}\\ \bm{\lambda}_{g_{1}}^{\top}&\mathbf{0}^{\top}\\ \mathbf{0}^{\top}&\bm{\lambda}_{g^{\prime}_{2}}^{\top}\end{pmatrix}=\begin{pmatrix}\bm{\lambda}^{*}_{1}&\bm{\lambda}^{*}_{g_{1}}&\mathbf{0}\\ \bm{\lambda}^{*}_{2}&\mathbf{0}&\bm{\lambda}^{*}_{g_{2}}\end{pmatrix}\begin{pmatrix}1&0&0\\ 0&1&\rho_{1,2}^{*}\\ 0&\rho_{1,2}^{*}&1\end{pmatrix}\begin{pmatrix}(\bm{\lambda}_{1}^{*})^{\top}&(\bm{\lambda}_{2}^{*})^{\top}\\ (\bm{\lambda}_{g_{1}}^{*})^{\top}&\mathbf{0}^{\top}\\ \mathbf{0}^{\top}&(\bm{\lambda}_{g_{2}}^{*})^{\top}\end{pmatrix},$ (G.3)
@@
$\begin{pmatrix}\bm{\lambda}_{1}&\bm{\lambda}_{g_{1}}&\mathbf{0}\\ \bm{\lambda}_{2,1}&\bm{\lambda}_{2,g_{1}}&\mathbf{0}\\ \bm{\lambda}_{2,2}&\mathbf{0}&\bm{\lambda}_{2,g^{\prime}_{2}}\end{pmatrix}\begin{pmatrix}1&0&0\\ 0&1&\rho_{1,2}\\ 0&\rho_{1,2}&1\end{pmatrix}\begin{pmatrix}\bm{\lambda}_{1}^{\top}&\bm{\lambda}_{2,1}^{\top}&\bm{\lambda}_{2,2}^{\top}\\ \bm{\lambda}_{g_{1}}^{\top}&\bm{\lambda}_{2,g_{1}}^{\top}&\mathbf{0}^{\top}\\ \mathbf{0}^{\top}&\mathbf{0}^{\top}&\bm{\lambda}_{2,g^{\prime}_{2}}^{\top}\end{pmatrix}$ (G.4)